OpenStudy (anonymous):

the perimeter of a rectangle is 10 ft. the length is 1 ft longer than the width .find the dimensions.writer a system of linear equations and solve the resulting systerm. let x be the length and y be the width. write the first equation.2x+2y= write the second equation. x=y+ what is the length?__ft what is the width?__ft

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Let x be the width of the rectangle. Then width=?

OpenStudy (anonymous):

sorry..length=?

OpenStudy (anonymous):

im so lost

OpenStudy (anonymous):

no you are not :) tell me what the length is ........

OpenStudy (anonymous):

right...a hint: the length is 1 ft longer than the width

OpenStudy (anonymous):

so if the width is x..then length =?

OpenStudy (anonymous):

i do not know

OpenStudy (anonymous):

i don't buy it..just read the above posts and gimme an answer.don't matter if its wrong

OpenStudy (anonymous):

1 and 5x

OpenStudy (anonymous):

what is 5x?

OpenStudy (anonymous):

length

OpenStudy (anonymous):

if x is width...5x is length. is that what you mean?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how can that be if the length is 1 ft longer than the width

OpenStudy (anonymous):

ok im dumb

OpenStudy (anonymous):

look..what is 2 in terms of 1??

OpenStudy (wolf1728):

A) 2x+2y= 10 B) x-y= 1 changing equation B x = y + 1 Substituting that into equation A) A) 2 *(y+1) + 2y = 10 2y + 2 + 2y =10 4y = 8 y = 2 feet x = Y +1 so x = 3 feet

OpenStudy (anonymous):

write the first equation.2x+2y=10 is right

OpenStudy (anonymous):

write the second equation. x=y+ is 1 right

OpenStudy (anonymous):

what is the length?__ft is 2feet what is the width?__ft is 3feet

OpenStudy (anonymous):

write the first equation.2x+2y=10 is right write the second equation. x=y+ is 1 right what is the length?__ft is 2feet what is the width?__ft is 3feet